OpenStudy (anonymous):

[5p-4]=[4p+31] those are absolute value bars.

OpenStudy (anonymous):

I can't remember what to do when there is an absolute value on each side of the equal sign,hmm... When there is just one, there is 2 equations that can be made. I'm guessing that when there is 2 we should be able to make 4 equations. \[1. \space 5p-4=4p+31\]\[2. \space 5p-4=-(4p+31)\]\[3. \space -(5p-4)=4p+31\]\[4. \space -(5p-4)=-(4p+31)\]

OpenStudy (anonymous):

Solve each equation, you might get 4 different answers for p

OpenStudy (anonymous):

I got (-3,35)

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

make sure you are distributing the negative sign to each term inside the ( )